# Can Large Language Models Beat Wall Street? Unveiling the Potential of AI in Stock Selection

### v1

@@ 5.2.3 GPT ranking
Building upon the insights gained from GPT-4’s rankings of MarketSenseAI signals, Figure 12 provides a visual representation of the quality of explanations accompanying MarketSenseAI’s ”buy” signals throughout the evaluation period. The figure depicts (bar plot) the frequency of ”buy” signals for each stock over the months. It also illustrates (scatter plot) the evaluative scores assigned by GPT-4, based on the depth and relevance of explanations behind each ”buy” recommendation. In this visualization, the scatter plot points are placed according to the average score given to each stock, allowing for an intuitive understanding of how the explanations for each stock were perceived in terms of quality and persuasiveness by the GPT-4 layer.
A key observation from Figure 12 is the higher scoring of technology and AI-related stocks, such as Nvidia, META, and Adobe. This trend is reflective of the market hype surrounding these sectors in 2023. The tendency of these stocks to receive higher scores not only demonstrates the relevance of MarketSenseAI’s ”buy” signals but also underscores the model’s capability to encapsulate and convey the prevailing market sentiment and potential in its explanations.
@@
6 Discussion
@@
The robust performance of MarketSenseAI within the S&P 100 — a heavily analyzed and traded stock universe — lays a solid foundation for potential outperformance in less scrutinized markets. In such markets, MarketSenseAI’s sophisticated analytical tools could offer even greater advantages, potentially uncovering opportunities that remain obscure to most market participants. This possibility paves the way for broader applications of MarketSenseAI across various financial assets, especially in markets traditionally lacking in-depth analysis.
However, it is vital to consider the limitations posed by the brief duration of the study. An extended evaluation period would offer a more thorough insight into MarketSenseAI’s effectiveness across diverse market conditions. This current limitation underscores the necessity for continued research and prolonged trials to confirm MarketSenseAI’s efficiency and flexibility throughout different market cycles and scenarios.
7 Conclusion
This work introduces MarketSenseAI, an innovative AI-driven framework for stock analysis and selection. Utilizing the advanced reasoning capabilities of GPT-4, MarketSenseAI effectively analyzes a diverse range of data, including company news, fundamentals, market dynamics, and macroeconomic information. In 2023, it demonstrated exceptional decision-making ability in both single stock recommendations and broader portfolio management within the S&P 100, yielding returns of up to 40% and surpassing market performance by 13%.
The findings underscore the significant benefits of integrating AI into financial analysis tools. MarketSenseAI enhances investment strategies with insights that are not only data-driven but also explainable. It offers a scalable and viable solution for retail and professional investors in single stock selection, leveraging them for more effective investment decisions.
Beyond the specific signals generated by MarketSenseAI, its individual components independently provide substantial value. They deliver a clear and concise picture of each factor affecting a stock, combining summarization capabilities with comprehensive data aggregation.
These results contribute to understanding the role of AI in finance, suggesting new possibilities for future advancements in this dynamic and evolving field. MarketSenseAI represents a step forward in the process of investment decision-making and shows potential in supporting the overall health and stability of the financial system.
@@ 7 Conclusion
Georgios Fatouros: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing – original draft & Visualization. Konstantinos Metaxas: Conceptualization, Methodology, Software, Validation, Investigation, Data curation, Writing - Review & Editing. John Soldatos: Funding acquisition, Writing – review & editing. Dimosthenis Kyriazis: Resources & Supervision.
@@
The authors declare their involvement in the development and planned commercialization of the presented framework (i.e., MarketSenseAI) as part of a fintech startup.
Funding
This work was supported by European Union’s funded Project HUMAINE [grant number 101120218].

### v2

@@ 5.2.3 GPT ranking
Building upon the insights gained from GPT-4’s rankings of MarketSenseAI signals, Figure 12 provides a visual representation of the quality of explanations accompanying MarketSenseAI’s ”buy” signals throughout the evaluation period. The figure depicts (bar plot) the frequency of ”buy” signals over the months for the stock with at least five ”buy” signals. It also illustrates (scatter plot) the evaluative scores assigned by GPT-4, based on the strength, depth and relevance of explanations behind each ”buy” recommendation. In this visualization, the scatter plot points are placed according to the average score given to each stock, allowing for an intuitive understanding of how the explanations for each stock were perceived in terms of quality and persuasiveness by the GPT-4 ranking layer.
@@
A key observation from Figure 12 is the significantly higher scoring of technology and AI-related stocks, such as Nvidia, Microsoft, and Amazon. This trend is reflective of the market hype surrounding these sectors during the evaluation period. The tendency of these stocks to receive higher scores not only demonstrates the relevance of MarketSenseAI’s ”buy” signals but also underscores the model’s capability to encapsulate and convey the prevailing market sentiment and potential in its explanations.
6 Conclusion
@@
The presented LLM-based system provides several key advantages over traditional financial tools, enhancing qualitative and quantitative data processing, mitigating biases, explaining its reasoning, and boosting scalability and efficiency. However, it’s crucial to acknowledge the study’s limitations, including the assumptions made about investment decision factors and GPT-4’s effectiveness in different markets. The limited duration of the study also highlights the need for extended research to fully understand MarketSenseAI’s effectiveness across various market conditions.
MarketSenseAI’s evaluation in the S&P 100, yielding up to 73% returns and surpassing market performance by 30%, showcases its potential in investment decision-making. In addition, beyond the specific signals generated by MarketSenseAI, its individual components independently provide substantial value. They deliver a clear and concise picture of each factor affecting a stock, combining summarization capabilities with comprehensive data aggregation.
These results contribute to understanding the role of LLMs in finance, suggesting new possibilities for future advancements in this dynamic and evolving field. MarketSenseAI represents a step forward in the process of investment decision-making and shows potential in supporting the overall health and stability of the financial system.
@@
George Fatouros: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing – original draft & Visualization. Kostas Metaxas: Conceptualization, Methodology, Software, Validation, Investigation, Data curation, Writing - Review & Editing. John Soldatos: Funding acquisition, Writing – review & editing. Dimosthenis Kyriazis: Resources & Supervision.
@@
The authors George Fatouros and Kostas Metaxas declare their involvement in the development and commercialization of the presented framework (i.e., MarketSenseAI) as a product of Alpha Tensor Technologies Ltd.